OpenStudy (nathalie0918):

A soccer ball is kicked toward the goal. The height of the ball is modeled by the function h(t) = -16t2 + 48t where t equals the time in seconds and h(t) represents the height of the ball at time t seconds. What is the axis of symmetry, and what does it represent?

jimthompson5910 (jim_thompson5910):

do you agree that the equation is in the form y = ax^2 + bx + c ?

OpenStudy (nathalie0918):

not really because I kind of new at this

jimthompson5910 (jim_thompson5910):

so you've never seen y = ax^2 + bx + c before? or \[\Large y = ax^2 + bx + c\] at all?

OpenStudy (nathalie0918):

no but I have to go but thanks for your help anyway